Grade 9Botany

a cistern can be filled separately by two pipes in 12 and 16 minutes sepratelyif both pipes are opened togehter when will the cistern be filled?

areeb anas
8 Years agoGrade 9

1 Answer

Arun
8 Years ago
Dear student
 
By first pipe...in 1 min = 1/12 part
By second pipe .....in 1 min = 1/16 part
Hence when both opened....in 1 min = 1/12 + 1/16
 = 4 + 3/48 = 7/48 part
Hence tank is filled in 48/7 min = 6. 857 min